OpenStudy (anonymous):

What is the approximate circumference of a circle with diameter fourteen centimeters? Use twenty-two sevenths to approximate pi.

OpenStudy (anonymous):

pi * d is the circumference of the circle where pi= 22/7=3.14 so the ans would be 3.14 * 14=43.9

OpenStudy (anonymous):

Wait... what... @Luis_Rivera First note that: \[ C=2\pi r=\pi d \]So, our circumference is: \[ 14\pi=14\left(\frac{22}{7}\right)=\frac{14}{7}\cdot 22=2\cdot22=44 \]

OpenStudy (anonymous):

(Well, never mind, now)

OpenStudy (anonymous):

Ok thanks i really needed this

OpenStudy (anonymous):

the answer is 44